Skip to Search

diagonal spacer

Information Technology Staff Meeting Notes
October 31, 2008

Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Jason Fleming, Winston Harris, Michael Jay, Paul McDonough, Chris Nicolich, Mark Sullivan, Evan Wack, Stephen Williams
  1. Matthew saw Grover at Sesame Street Live!
  2. HR - minimum wage to increase to $7.21 in January.
  3. Xerox contract expires in November. The new vendor on the campus contract will include non-affiliates. We will save money on printing.
  4. Branch manager meeting - they are actively considering opening the Libraries 24 hours per day, five days per week. They are looking at SG to help with funding.
  5. MMM - talk of walk-up computers for walk-ins to access only the catalog.
Review Minutes of 10/17/08
  1. Web #3 - Matthew's IGoogle widget really works. It's the mobile cell that works only in theory.
Ongoing IT Updates
Equipment Received/Installed/Repaired
  1. Orders - MSL battery saga - still ongoing.
  2. Room 1A - installed O2K7. Need to uninstall and put O2K3 back on with the compatibility pack.
  3. Power strips and network cables are being stolen in MSL. There are 89 tables that can use power domes - Facilities will bolt them down.
  4. Cynthia is having trouble with Bonnie Smith's Tungsten Handheld device not synching groups and contacts.
Support Issues
  1. SMS client has been installed on most workstations. There are 46 that won't install. Amy will create a ticket for the Hardware Support team to install those by hand.
  2. Steady State testing - problem with the public image. Right now it is installed in the media area and high-end machines in West as well as on the Bloomberg.
  3. New DHCP image created this morning. No working image for West. Problems with the updates.
  1. Erich agreed to use Tivoli with random checks. Mark says that DLC is actually holding off now on Tivoli. They are checking with FDA at FCLA to see if they are capable. They have no hard limit. We started pushing to them yesterday.
  2. SAN migration - we still have to move Library West and Projects as well as Administration. Logan will do it upon return from vacation.
  3. USB issue - non-administrators unable to install. We are still looking for a solution for this.
  4. XP service packs - Amy has a solution and is working on it.
  5. CJK - nothing more for Logan to do, really. It is now a matter of how they want it set up. It can be done by the workstation support unit. With Steady State, it will reboot each time. Discussion on set up options. We will get with David Hickey and public services to get a decision on the set up. We need to move the machine to the back, instead of it being right out there in the middle of the floor.
  6. Campus AD planning - waiting for UFAD to modify the file for us. We hope to migrate workstations during the first week of December. Users will have to log into Outlook during the week gap until email is migrated.
  7. Disaster Prep - Achilles Program - a campus requirement - need to have it ready by the end of November.
Workstation Support Unit
  1. West re-imaging is on hold.
  2. Inventory will be worked on soon.
  3. Cable ties are on order to clean up tangled areas.
  4. Laptop batteries - reimbursing by SPAN.
  1. Machine in DLC - ALEPH is not working on it. Same situation as in LAC. Jason is working on it.
  2. CSUL will decide whether we will use ALEPH 19 or ALEPH 20.
Software Projects Status
  1. VIVO/Gator Scholar - struggled this week, but finally got it successfully installed and loaded Val and Sara's information onto it. Stephen is documenting the process.
  2. ARL stats - Michael is working on the tables.
  3. DLC Tracking - problem with the server this morning. It had been pushed updates and was waiting for a restart - causing all services to stop. The restart needs to be automated. The problem is with SQL Express. The settings need to be changed for a reboot.
  4. The Backup Server has a drive problem. After the reboot, the drive problems have increased. Need to move it to a different server.
  5. UFDC - we now have a couple of our collections that are so big, Greenstone can't index them. Mark is switching to a different index engine.
  6. DLC is now mapping between ALEPH, MARC and ENDECA (Mango), so that digital resources appear in the Library Catalog.
  7. Patron Updates - we had two requests for changes. Chris has finished one and the second is close to completion. This change will allow students auditing a class to be able to check out library books.
  8. Maps - Hesham made changes - it now will not work outside out domain. He is out of the country for awhile. We will try to get it to work if possible, but it's not top priority.
  9. MSL public computers now have the availability tool. The link has been put up on the Web.
Students and Department
  1. Erasmo is no longer employed in Library IT.
  2. Two new work study students have been hired: Caitlyn Mertz and Lorraine Wowo.
  3. With newly employed students, please remember to keep PAULA up-to-date.
IT Security Management
  1. Meeting held.


University of Florida Home Page